# Protein Conformation Generation via Force-Guided $\mathrm{SE}(3)$ Diffusion Models

### v1

@@ Abstract
The conformational landscape of proteins is crucial to understanding their functionality in complex biological processes. Traditional physics-based computational methods, such as molecular dynamics (MD) simulations, suffer from rare event sampling and long equilibration time problems, hindering their applications in general protein systems. Recently, deep generative modeling techniques, especially diffusion models, have been employed to generate novel protein conformations. However, existing score-based diffusion methods cannot properly incorporate important physical prior knowledge to guide the generation process, causing large deviations in the sampled protein conformations from the equilibrium distribution. In this paper, to overcome these limitations, we propose a force-guided $\mathrm{SE}(3)$ diffusion model, ConfDiff, for protein conformation generation. By incorporating a force-guided network with a mixture of data-based score models, ConfDiff can can generate protein conformations with rich diversity while preserving high fidelity. Experiments on a variety of protein conformation prediction tasks, including 12 fast-folding proteins and the Bovine Pancreatic Trypsin Inhibitor (BPTI), demonstrate that our method surpasses the state-of-the-art method.
@@ 3.1 Protein Backbone Diffusion on SE(3)
[…]
@@ 3.2 Classifier-free Guidance for Diffusion Sampling
$=$ $\gamma\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t},c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$
@@ 4.1 Sequence-Conditional Diffusion on $\mathrm{SE}(3)$
4.2 Intermediate Energy Guided Diffusion
@@ 5 Experiments
We evaluate the effectiveness of ConfDiff with different guidance strategies on protein conformation generation tasks. In brief, we demonstrate that 1) classifier-free guidance balances the trade-off between sample quality and diversity; 2) energy and force guidance effectively guide the model to sample conformations with lower energy; 3) combining classifier-free and physical guidance strategies, ConfDiff outperforms current state-of-the-art models by generating high-quality and diverse samples more truthful to the underlying Boltzmann distribution.
@@ 5.1 Experimental Setup
Datasets. We apply ConfDiff on two protein conformation benchmarks for a comprehensive evaluation: 1) fast-folding proteins comprise MD simulation data for 12 small proteins with fast folding-unfolding dynamics (Lindorff-Larsen et al., ), 2) bovine pancreatic trypsin inhibitor (BPTI) contains MD simulation data for BPTI exhibiting five metastable states (Shaw et al., 2010). See Appendix C.1 for more details.
@@
PwD Rg TIC TIC-2D ($\mathrm{\SIUnitSymbolAngstrom}$) ($\mathrm{\SIUnitSymbolAngstrom}$)
@@
Metrics: we evaluate generated conformations as to their validity, precision, diversity and the similarity of distributions when comparing conformation ensembles. We summarize key metrics here and include details in Appendix C.4:
@@ 5.2 Conformation Sampling with Force Guidance
As an initial proof-of-concept, we first show that proposed force guidance can improve the sampling capability of ConfDiff by generating protein conformations with reduced energy (i.e. improved stability). We select WW-domain, one of the fast-folding proteins from Lindorff-Larsen et al. as an example and generate conformations using ConfDiff at varying degrees of force guidance ($\eta$) and sequence conditions ($\gamma$). As shown in Figure 2, without force guidance ($\eta=0$), models with weaker sequence condition can generate samples with higher diversity (RMSF), however, also with higher energy, showing the quality-diversity compromise. After integrating force guidance, ConfDiff mitigates such effect and are able to generates samples with lower energy while maintaining similar RMSF levels. This result shows that force guidance can improve conformation stability without drastically decease the diversity. This favorable effect suggests integrating physical information might provide fine-grained guidance to sample locally optimized structures.
Figure 2: Energy (left) and diversity (right) of sampled conformations for WW Domain with various levels of force guidance ($\eta$) and sequence conditions ($\gamma$). Models with weaker sequence condition generate more diverse samples and force guidance improves conformation stability without drastically decease the diversity.
Figure 3: Sample distribution over first two TIC components for WW Domain. Experimental structure in the folded state is shown in color; 5 random samples from reference MD are shown as grey.
@@ 5.3 Distribution Prediction for Fast-folding Proteins
[…]
@@ 5.4 Metastable States Prediction for BPTI
Previous MD simulation studies on BPTI has recovered 5 five kinetic clusters representing metastable states near its native folded state (Shaw et al., 2010). For this benchmark, we evaluate the quality and efficiency for models to recover these 5 states. Compared with fast-folding proteins, this benchmark requires model to generate different conformations while maintaining the correct folding structure.
We measure model precision by the lowest sample RMSD to cluster centers. For comparison, we report the best RMSD averaged over 5 clusters (RMSDavg) and the RMSD to cluster 3 (RMSDcls3), the cluster that is the most difficult to sample. As shown in Table 2, ConfDiff with force guidance performs the best on both metrics, suggesting better capability on predicting distinct metastable states.
Table 2: Recovery of 5 metastable states for BPTI. RMSDavg denotes the best RMSD score, averaged over 5 states. RMSDcl3 refers to the best RMSD to the remote Cluster 3, which is the most difficult to sample. (Unit: $\mathrm{\SIUnitSymbolAngstrom}$.)
@@ B.1 Diffusion on $\mathrm{SE(3)}$
$\mathrm{d}\mathbf{x}_{t}=\left[-\frac{1}{2}\beta_{t}\mathbf{P}\mathbf{T}_{t},0\right]\mathrm{d}t+\left[\sqrt{\beta_{t}}\mathbf{P}\mathrm{d}\mathbf{w}_{t},\sqrt{\frac{d}{dt}\sigma^{2}_{t}}\mathrm{d}\mathbf{w}^{\mathrm{SO}(3)}_{t}\right],\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$
@@
$\mathrm{d}\mathbf{x}_{t}=\left[\mathbf{P}\left(-\frac{1}{2}\beta_{t}\mathbf{T}_{t}-\beta_{t}\nabla\log q_{t}(\mathbf{T}_{t})\right),-\frac{d}{dt}\sigma^{2}_{t}\nabla\log q_{t}(\mathbf{R}_{t})\right]\mathrm{d}t+\left[\sqrt{\beta_{t}}\mathbf{P}\mathrm{d}\bar{\mathbf{w}}_{t},\sqrt{\frac{d}{dt}\sigma^{2}_{t}}\mathrm{d}\bar{\mathbf{w}}^{\mathrm{SO}(3)}_{t}\right].$
@@ B.2 Interpolation with Simulation Data
According to the derivation in Appendix A, $\nabla_{\mathrm{x}_{0}}\mathcal{E}_{0}(\mathrm{x}_{0})$ has the ground truth value at the boundary $t\in\{0,1\}$ and we observe that both score functions $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ and $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ exhibit large variance at $t=0$ (due to small noise level at small $t$). Thus, we propose an interpolation strategy for the intermediate force function like in (Máté & Fleuret, 2023):
@@
Nevertheless, throughout the inference process, despite the ability to derive a predicted $\widehat{\mathbf{x}}_{0}$ at every step for the computation of $\nabla_{\widehat{\mathbf{x}}_{0}}\mathcal{E}_{0}(\widehat{\mathbf{x}}_{0})$, the utilization of OpenMM for energy calculation is highly time-consuming, particularly in estimating energy in lengthy protein chains. Thus, we additionally train a network $\widetilde{g}_{\nu}(\mathbf{x}_{t},t)$ to estimate the ground truth force field $\nabla_{\mathbf{x}_{0}}\mathcal{E}_{0}(\mathbf{x}_{0})$ at the initial moment by MSE loss function $\mathcal{L}_{t=0}(\nu)$, resulting in the transformation of $h_{\psi}(\mathbf{x}_{t},t)$ into
@@ B.3 Implementation
The single layer of ConfDiff consists of invariant point attention layer (Jumper et al., 2021), transformer and other structure modules similar to  Yim et al. (2023b). In the unconditional diffusion model, the single representation $s_{0}$ is initialized as a concatenation of the position encoding of residues and sinusoidal time embedding. Meanwhile, the pair representation $z_{0}$ is initialized as a concatenation of the relative residue position encoding, sinusoidal time embedding, and relative sequence distances. In the conditional model, the node embedding of ESMfold Lin et al. (2022) with corresponding sequences is additionally concatenated to the single representation $s_{0}$. In addition, we apply the same noise-injection schedule as the translation term in Eq. (1) to the ESMFold node embeddings before feeding to the neural networks. This noise injection step further increases sampling diversity. In both the intermediate force prediction network ${g}_{\nu}(\mathbf{x}_{t},t)$ and the force-field prediction network $\widetilde{g}_{\nu}(\mathbf{x}_{t},t)$, the features are initialized as the output single and pair representation of the conditional model. These features are then used to predict the intermediate force, which is derived from the last three dimensions of the newly predicted rigids. The intermediate energy prediction network $f_{\phi}(\mathbf{x}_{t},t)$ also takes the output of the conditional model as its input and a linear layer is added at the end that transforms the single representation into a scalar value, thereby aiming to predict energy. Detailed hyperparameters are listed in S1.
B.4 Hyperparameters
@@ C.1 Datasets
Fast folding proteins. The simulation details of fast-folding proteins are described in (Lindorff-Larsen et al., ) and the trajectories are obtained from the authors. The dataset contains 12 small proteins range from 10 to 80 residues that exhibit multiple fast folding-unfolding events within the millisecond-scale all-atom simulations. Atom coordinates are saved at 200 $\mathrm{ps}$ intervals and 520K $\sim$ 14.7M aggregated frames are collected for each protein. $\alpha$-carbon atom coordinates are used for analysis. For sampling, we generate 1000 conformations for each proteins with guidance strength $\eta=1.3$ for both force and energy guidance, and the sequence condition $\gamma$ sampled between $0.5\sim 1.0$.
@@ C.2 Time-lagged Independent Analysis (TICA)
For fast-folding proteins and BPTI, we use pyemma (Scherer et al., 2015) for featurizing protein structures (as pairwise distances between $\alpha$-carbons), and use Deeptime (Hoffmann et al., 2021) for time-lagged independent analysis (TICA) and projecting the protein conformations into two slowest components for distribution comparison (Naritomi & Fuchigami, 2013; Pérez-Hernández et al., 2013).

### v2

@@ Abstract
The conformational landscape of proteins is crucial to understanding their functionality in complex biological processes. Traditional physics-based computational methods, such as molecular dynamics (MD) simulations, suffer from rare event sampling and long equilibration time problems, hindering their applications in general protein systems. Recently, deep generative modeling techniques, especially diffusion models, have been employed to generate novel protein conformations. However, existing score-based diffusion methods cannot properly incorporate important physical prior knowledge to guide the generation process, causing large deviations in the sampled protein conformations from the equilibrium distribution. In this paper, to overcome these limitations, we propose a force-guided $\mathrm{SE}(3)$ diffusion model, ConfDiff, for protein conformation generation. By incorporating a force-guided network with a mixture of data-based score models, ConfDiff can generate protein conformations with rich diversity while preserving high fidelity. Experiments on a variety of protein conformation prediction tasks, including 12 fast-folding proteins and the Bovine Pancreatic Trypsin Inhibitor (BPTI), demonstrate that our method surpasses the state-of-the-art method.
@@ 3.1 Protein Backbone Diffusion on SE(3)
We adopt the protein backbone representation from AlphaFold (Jumper et al., 2021): for a protein with $N$ amino acid residues, its backbone atomic coordinates can be parameterized by a collection of $N$ orientation preserving rigid transformations (i.e., frames) to the local [N,Cα,C,O] backbone atoms in each residue. We collectively denote the positions of all $N$ frames by $\mathbf{x}_{0}=[\mathbf{T}_{0},\mathbf{R}_{0}]\in\mathrm{SE}(3)^{N}$, where $\mathbf{T}_{0}\in\mathbb{R}^{3N}$ and $\mathbf{R}_{0}\in\mathrm{SO(3)}^{N}$ denote the corresponding translation and rotation operations, respectively. With an additional backbone torsion angle $\psi$ describing the rotation of the oxygen atom around the C–Cα bond within each residue, we can reconstruct the protein backbone structure from the frame representations.
@@ 3.2 Classifier-free Guidance for Diffusion Sampling
$=$ $\gamma\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$
@@ 4.1 Sequence-Conditional Diffusion on $\mathrm{SE}(3)$
4.2 Intermediate Energy-Guided Diffusion
@@ 5 Experiments
We evaluate the effectiveness of ConfDiff with different guidance strategies on protein conformation generation tasks. In brief, we demonstrate that 1) classifier-free guidance balances the trade-off between sample quality and diversity; 2) energy and force guidance effectively guide the model to sample conformations with lower energy; 3) combining classifier-free and physical guidance strategies, ConfDiff outperforms current state-of-the-art models by generating high-quality and diverse samples more truthful to the underlying Boltzmann distribution. Code and checkpoints are open-sourced at https://github.com/bytedance/ConfDiff.
@@ 5.1 Experimental Setup
Datasets. We apply ConfDiff on two protein conformation benchmarks for a comprehensive evaluation: 1) fast-folding proteins comprise MD simulation data for 12 small proteins with fast folding-unfolding dynamics (Lindorff-Larsen et al., 2021), 2) bovine pancreatic trypsin inhibitor (BPTI) contains MD simulation data for BPTI exhibiting five metastable states (Shaw et al., 2010). See Appendix C.1 for more details.
@@
PwD Rg TIC TIC-2D ($\uparrow$) ($\downarrow$) ($\mathrm{\SIUnitSymbolAngstrom}$)
@@
Metrics. We evaluate generated conformations as to their validity, precision, diversity and the similarity of distributions when comparing conformation ensembles. We summarize key metrics here and include details in Appendix C.4:
@@ 5.2 Conformation Sampling with Force Guidance
As an initial proof-of-concept, we first show that proposed force guidance can improve the sampling capability of ConfDiff by generating protein conformations with reduced energy (i.e. improved stability). We select WW-domain, one of the fast-folding proteins from Lindorff-Larsen et al. (2021) as an example and generate conformations using ConfDiff at varying degrees of force guidance ($\eta$) and sequence condition ($\gamma$). As shown in Figure 2, without force guidance ($\eta=0$), models with weaker sequence condition can generate samples with higher diversity (RMSF), however, also with higher energy, showing the quality-diversity compromise. After integrating force guidance, ConfDiff mitigates such effect and are able to generates samples with lower energy while maintaining similar RMSF levels. This result shows that force guidance can improve conformation stability without drastically decease the diversity. This favorable effect suggests integrating physical information might provide fine-grained guidance to sample locally optimized structures.
Figure 2: Energy (left) and diversity (right) of sampled conformations for WW-domain with various levels of force guidance ($\eta$) and sequence condition ($\gamma$). Models with weaker sequence condition generate more diverse samples and force guidance improves conformation stability without drastically decease the diversity.
Figure 3: Sample distribution over the first two TIC components for WW-domain. Ref (N=1000) shows 1000 random samples from the reference MD simulation. The illustration in the lower left shows the experimental structure in the folded state (in color) and 5 random samples from the reference (in grey).
@@ 5.3 Distribution Prediction for Fast-Folding Proteins
This dataset consists of 12 short proteins whose folding and unfolding dynamics have been examined in long-time all-atom MD simulations (Lindorff-Larsen et al., 2021). In this experiment, we evaluate the capability of models to recover the conformation distributions observed in the simulation. This benchmark requires the model to not only predict the stable folded structures, but also physically viable unfolded or partially unfolded structures. To assess distribution similarity, we evaluate JS-distances between the distributions of generated samples and ground truth MD samples, along with the validity score and RMSF for diversity. In addition, we report the root-mean-square-error on the predicted contact rate between residues RMSEcontact to reflect the model accuracy in flexible regions (Janson et al., 2023).
@@ 5.4 Metastable States Prediction for BPTI
Previous MD simulation studies on BPTI have recovered five kinetic clusters, representing metastable states near its native folded state (Shaw et al., 2010). For this benchmark, we evaluate the quality and efficiency for models to recover these five states. Compared with fast-folding proteins, this benchmark requires model to generate different conformations while maintaining the correct folding structure.
We measure model precision by the lowest sample RMSD to cluster centers. For comparison, we report the best RMSD averaged over five clusters (RMSDavg) and the RMSD to cluster 3 (RMSDcls3), the cluster that is the most difficult to sample. As shown in Table 2, ConfDiff with force guidance performs the best on both metrics, suggesting better capability on predicting distinct metastable states.
Table 2: Recovery of five metastable states for BPTI. RMSDavg denotes the best RMSD score, averaged over five states. RMSDcls3 refers to the best RMSD to the remote cluster 3, which is the most difficult to sample. (Unit: $\mathrm{\SIUnitSymbolAngstrom}$.)
@@ B.1 Diffusion on $\mathrm{SE(3)}$
$\mathrm{d}\mathbf{x}_{t}=\left[-\frac{1}{2}\beta_{t}\mathbf{P}\mathbf{T}_{t},0\right]\mathrm{d}t+\left[\sqrt{\beta_{t}}\mathbf{P}\mathrm{d}\mathbf{w}_{t},\sqrt{\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}_{t}}\mathrm{d}\mathbf{w}^{\mathrm{SO}(3)}_{t}\right],\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$
@@
$\mathrm{d}\mathbf{x}_{t}=\left[\mathbf{P}\left(-\frac{1}{2}\beta_{t}\mathbf{T}_{t}-\beta_{t}\nabla\log q_{t}(\mathbf{T}_{t})\right),-\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}_{t}\nabla\log q_{t}(\mathbf{R}_{t})\right]\mathrm{d}t+\left[\sqrt{\beta_{t}}\mathbf{P}\mathrm{d}\bar{\mathbf{w}}_{t},\sqrt{\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}_{t}}\mathrm{d}\bar{\mathbf{w}}^{\mathrm{SO}(3)}_{t}\right].$
@@ B.2 Interpolation with Simulation Data
According to the derivation in Appendix A, $\nabla_{\mathrm{x}_{0}}\mathcal{E}_{0}(\mathrm{x}_{0})$ has the ground truth value at the boundary $t\in\{0,1\}$ and we observe that both score functions $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ and $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ exhibit large variance at $t=0$ (due to small noise level at small $t$). Thus, we propose an interpolation strategy for the intermediate force function like in (Máté & Fleuret, 2023; Wang et al., 2024):
@@
Nevertheless, throughout the inference process, despite the ability to derive a predicted $\widehat{\mathbf{x}}_{0}$ at every step for the computation of $\nabla_{\widehat{\mathbf{x}}_{0}}\mathcal{E}_{0}(\widehat{\mathbf{x}}_{0})$, the utilization of OpenMM for energy calculation is highly time-consuming, particularly in estimating energy in lengthy protein chains. To further demonstrate the computational cost of potential energy evaluation, we tabulate the total time consumption of once energy/force calculation using openMM (the cost of side-chain packing using faspr is excluded) for the fast-folding proteins. Due to the 1000 steps in the inference diffusion process, the time required to generate a single protein conformation is multiplied by 1000.
Table S1: The time consumption of once energy/force calculation using openMM.
1FME 2F4K 2JOF 2WAV A3D CLN025
Length 28 35 20 47 73 10
CPU 3.6±0.6 s 4.2±1.8 s 1.4±0.3 s 8.3±3.3 s 19.4±5.3 s 0.9±0.2 s
GPU 1.7±0.2 s 1.7±0.5 s 1.6±0.5 s 1.9±0.6 s 2.2±0.5 s 1.6±0.5 s
GTT NTL9 NuG2 PRB UVF lambda
Length 35 39 56 47 52 80
CPU 5.0±2.1 s 6.8±1.4 s 12.1±5.5 s 8.7±1.6 s 12.6±2.5 s 26.6±16.4 s
GPU 1.8±0.5 s 1.8±0.6 s 2.0±0.5 s 1.8±0.5 s 2.0±0.5 s 2.0±0.5 s
Considering that the reverse diffusion process is discretized into 1000 steps, the time expenditure is not feasible to incorporate the ground-truth MD force field in each reverse sampling step, necessitating the employment of a surrogate model. Thus, we additionally train a network $\widetilde{g}_{\nu}(\mathbf{x}_{t},t)$ to estimate the ground truth force field $\nabla_{\mathbf{x}_{0}}\mathcal{E}_{0}(\mathbf{x}_{0})$ at the initial moment by MSE loss function $\mathcal{L}_{t=0}(\nu)$, resulting in the transformation of $h_{\psi}(\mathbf{x}_{t},t)$ into
@@ B.3 Implementation
The single layer of ConfDiff consists of invariant point attention layer (Jumper et al., 2021), transformer and other structure modules similar to  Yim et al. (2023b). In the unconditional diffusion model, the single representation $s_{0}$ is initialized as a concatenation of the position encoding of residues and sinusoidal time embedding. Meanwhile, the pair representation $z_{0}$ is initialized as a concatenation of the relative residue position encoding, sinusoidal time embedding, and relative sequence distances. In the conditional model, the node embedding of ESMfold Lin et al. (2022) with corresponding sequences is additionally concatenated to the single representation $s_{0}$. In addition, we apply the same noise-injection schedule as the translation term in Eq. (1) to the ESMFold node embeddings before feeding to the neural networks. This noise injection step further increases sampling diversity. In both the intermediate force prediction network ${g}_{\nu}(\mathbf{x}_{t},t)$ and the force-field prediction network $\widetilde{g}_{\nu}(\mathbf{x}_{t},t)$, the features are initialized as the output single and pair representation of the conditional model. These features are then used to predict the intermediate force, which is derived from the last three dimensions of the newly predicted rigids. The intermediate energy prediction network $f_{\phi}(\mathbf{x}_{t},t)$ also takes the output of the conditional model as its input and a linear layer is added at the end that transforms the single representation into a scalar value, thereby aiming to predict energy. Detailed hyperparameters are listed in S3.
B.4 Computational Cost
Our ConfDiff-base model has similar computational complexity to other diffusion-based conformation generation models, e.g., FrameDiff (Yim et al., 2023b) and Str2Str (Lu et al., 2024). The introduction of a guidance network will not significantly increase computational cost. Specifically, the energy/force guidance modules only introduce a small computational overhead during inference through a lightweight energy/force prediction head, which is based on the pretrained ConfDiff-base model to accomplish guidance.
We made a comparative study of the time and GPU memory requirement for generating protein conformations across different sequence lengths. We used a single NVIDIA-V100 GPU to benchmark model performance. The batch size is 16, and the number of reverse sampling steps is 1000.
Table S2: The time and memory requirement for generating protein conformations across different sequence lengths.
ConfDiff-base ConfDiff-Energy ConfDiff-Force
len (protein) Time (sec) Memory (MB) Time (sec) Memory (MB) Time (sec) Memory (MB)
58 (BPT) 71 1838 110 3142 104 1916
118 (1JM4.B) 167 3932 256 9722 206 4010
238 (2LAO.A) 578 12402 888 29886 709 12480
Notably, the energy guidance policy almost doubles the memory cost compared to ConfDiff-base, since the derivative of energy is required to compute the force guidance term. Meanwhile, force guidance only consumes negligibly more computational resources compared to ConfDiff-base, which is another significant advantage of the force guidance strategy versus the energy guidance strategy.
B.5 Hyperparameters
@@ C.1 Datasets
Fast-folding proteins. The simulation details of fast-folding proteins are described in (Lindorff-Larsen et al., 2021) and the trajectories are obtained from the authors. The dataset contains 12 small proteins range from 10 to 80 residues that exhibit multiple fast folding-unfolding events within the millisecond-scale all-atom simulations. Atom coordinates are saved at 200 $\mathrm{ps}$ intervals and 520K $\sim$ 14.7M aggregated frames are collected for each protein. $\alpha$-carbon atom coordinates are used for analysis. For sampling, we generate 1000 conformations for each proteins with guidance strength $\eta=1.3$ for both force and energy guidance, and the sequence condition $\gamma$ sampled between $0.5\sim 1.0$.
@@ C.2 Time-Lagged Independent Analysis (TICA)
For fast-folding proteins and BPTI, we use pyemma (Scherer et al., 2015) for featurizing protein structures (as pairwise distances between $\alpha$-carbons), and use Deeptime (Hoffmann et al., 2021) for time-lagged independent analysis (TICA) and projecting the protein conformations into the two slowest components for distribution comparison (Naritomi & Fuchigami, 2013; Pérez-Hernández et al., 2013).